# Visual Riddles: a Commonsense and World Knowledge Challenge for Large Vision and Language Models

### v1

@@ A.4 Prompts for Different Models
  1. 1.
    Multiple-Choice VQA: A specific prompt is used for generating distractors when there are insufficient incorrect answers using GPT-4. Additionally, for multiple-choice VQA, we evaluate the models’ performance across three settings: (1) given only the image, the question, and five possible answers (with only one correct answer), (2) given the image, the question, the possible answers, and a hint, and (3) given the image, the question, the possible answers, and an attribution. The prompts structure is outlined in Table 9.
  2. 2.
    VQA Automatic Evaluation: To find the best judge (auto-rater), we evaluate the models in two scenarios - reference-free and reference-based. The prompts structure is outlined in Table 10.
  3. 3.
    Auto-Rater for Open-Ended VQA and Ablation Study with Modified Images: Using the best Auto-Rater we evaluate all models automatic ratings on the Open-Ended VQA task. We also use the same prompt in our analysis we perform ablation study to explores whether models base their answers solely on text or consider visual clues. The prompts structure is outlined in Table 10.
@@ A.6 Open-Ended VQA: Humans and Models Answers Annotation Evaluation Guidelines
In order to evaluate how well models and humans answered the visual riddle, we used MTurk platform, paying 18\$ per hour. We contacted workers and invited them to a qualification round that began with a review of task guidelines and included solving five riddles of varying difficulty. Following an assessment of their responses and providing personalized feedback, only those with a strong understanding of the task moved on to the actual annotation phase. Of 9 candidates, 6 proceeded to the actual annotation phase.

### v2

@@ A.3 Categories and Difficulty Level: Breakdown and Results
Figure 10: Aggregated results for model and human performance across categories (top) and difficulty levels (middle), along with visual riddle examples illustrating varying difficulty levels (bottom). The top section shows the average open-ended VQA accuracy for models and humans by category, with a correlation of -0.12 between their performances. The middle section presents the average open-ended VQA accuracy for models and humans by difficulty level, showing a correlation of 0.94.
@@
To provide a clearer understanding of the performance across difficulty levels, we included aggregated results for open-ended VQA accuracy of both humans and models, categorized by different categories and difficulty index levels (see Figure 10). The analysis shows that certain categories, such as Temporal Principles, Object Counting, and Biological Principles, pose greater challenges for models, while Geographic Knowledge and Physical Principles are more difficult for humans. Interestingly, there is no significant correlation between the difficulty categories for humans and models, with a correlation coefficient of -0.12. For difficulty levels, while human performance consistently surpasses models, both show a decline as difficulty increases, with a strong correlation of 0.94 (see top and middle parts of Fig. 9). To further illustrate, examples of images corresponding to different difficulty levels are provided in Fig. 10 (bottom).
@@ A.4 Prompts for Different Models
Multiple-Choice VQA:
A specific prompt is used for generating distractors when there are insufficient incorrect answers using GPT-4. Additionally, for multiple-choice VQA, we evaluate the models’ performance across three settings: (1) given only the image, the question, and five possible answers (with only one correct answer), (2) given the image, the question, the possible answers, and a hint, and (3) given the image, the question, the possible answers, and an attribution. The prompts structure is outlined in Table 9.
VQA Automatic Evaluation:
To find the best judge (auto-rater), we evaluate the models in two scenarios - reference-free and reference-based. The prompts structure is outlined in Table 10.
Auto-Rater for Open-Ended VQA and Ablation Study with Modified Images:
Using the best Auto-Rater we evaluate all models automatic ratings on the Open-Ended VQA task. We also use the same prompt in our analysis we perform ablation study to explores whether models base their answers solely on text or consider visual clues. The prompts structure is outlined in Table 10.
@@ A.6 Open-Ended VQA: Humans and Models Answers Annotation Evaluation Guidelines
In order to evaluate how well models and humans answered the visual riddle, we used MTurk platform, paying 18\$ per hour. We contacted workers and invited them to a qualification round that began with a review of task guidelines and included solving five riddles of varying difficulty. Following an assessment of their responses and providing personalized feedback, only those with a strong understanding of the task moved on to the actual annotation phase. Of 9 candidates, 6 proceeded to the actual annotation phase. To ensure greater cultural diversity in the benchmark, all AMT annotators were selected from countries different from those of the visual riddles creators.